# Modeling the Siege of Syracuse: Resources, strategy, and collapse

### v1

@@ Abstract
The Siege of Syracuse (214–212 BC) was a decisive event in the Second Punic War, leading to the city’s fall to Rome despite its formidable defenses, including the war machines devised by Archimedes. In this work, we propose a mathematical model to describe the dynamics of the siege, incorporating the depletion of resources, the decline of the Syracusan population, and the persistence of the Roman army. Our analysis reveals the existence of a critical threshold $\lambda_{c}$, which determines the outcome of the siege. This threshold marks a phase transition: if the effectiveness of Syracuse’s defenses, represented by $\lambda$, had exceeded $\lambda_{c}$, the city could have withstood the Roman assault. However, since history records Syracuse’s fall, we conclude that $\lambda<\lambda_{c}$. This result provides a quantitative framework to understand the inevitability of the city’s conquest and demonstrates how mathematical modeling can offer new insights into historical military conflicts. We also explore different scenarios and assess the impact of key factors such as siege duration, supply constraints, and defensive capabilities. The results provide insight into how strategic elements influenced the eventual fall of Syracuse and demonstrate the applicability of mathematical modeling in historical military analysis.
@@ 1 Introduction
The siege of Syracuse (214–212 BC) was a pivotal event during the Second Punic War, in which the Roman Republic besieged and ultimately conquered the powerful Greek city-state of Syracuse, located in Sicily. This siege marked a crucial moment in Rome’s expansionist ambitions and solidified its control over the Mediterranean. Syracuse, previously an ally of Rome under King Hiero II, shifted its allegiance to Carthage after Hiero’s death, prompting a Roman military response under the command of Marcus Claudius Marcellus [1].
@@
The fall of Syracuse can be understood in terms of a critical threshold in the effectiveness of its defenses. In the mathematical model developed in this work, we identify a critical value $\lambda_{c}$ that marks a transition between two possible outcomes: the survival of Syracuse or its eventual conquest by Rome. This behavior resembles a phase transition in physics, where small variations in parameters lead to qualitatively different macroscopic states. If the defensive effectiveness $\lambda$ had exceeded the critical threshold $\lambda_{c}$, Syracuse might have been able to resist the Roman siege indefinitely. However, since historical records indicate the city’s fall, we conclude that $\lambda<\lambda_{c}$, confirming that the defenses, despite their ingenuity—including the war machines attributed to Archimedes—were ultimately insufficient to prevent Roman victory. This approach provides a novel perspective on the siege, connecting historical events to universal principles of dynamical systems and critical phenomena.
Thus, in this work, we develop a system of three differential equations to model the interactions between available resources, the besieged population, and the Roman attacking force. These equations capture the depletion of resources over time, the decline in the defending population due to starvation and combat, and the attrition of the Roman army. By solving these equations numerically, we explore different scenarios and assess the conditions under which Syracuse ultimately fell.
@@ 2 Mathematical model
To analyze the dynamics of the Siege of Syracuse, we develop a system of three differential equations representing the interaction between available resources, the besieged population, and the attacking Roman forces. The model captures key factors such as resource depletion, attrition due to starvation and combat, and the effectiveness of defensive and offensive strategies.
@@
The first equation models the decline of resources due to consumption by the population and natural degradation over time. The second equation captures population losses due to natural causes, starvation, and direct military confrontations, but it presents a term that models the survivance due to limited resources. The third equation represents the attrition of the Roman army caused by resistance from the defenders, and Roman army reinforcement.
@@ 3 Results
From Eq. (3), we find two steady-state solutions: $A=0$ and $P=\mu/\lambda$. Using the steady-state value of $R$, Eq. (2) yields:
$P=\frac{\delta\beta}{\alpha(\delta+\kappa A)}.$ (5)
For the case where $A=0$, we obtain $P=\frac{\delta\beta}{\alpha\delta}$, leading to the first stationary solution:
$(R,A,P)=\left(\frac{\beta}{\alpha},0,\frac{\delta\beta}{\alpha\delta}\right).$ (6)
On the other hand, if $A\neq 0$, we find $P=\frac{\mu}{\lambda}$. This result, together with Eq. (5), leads to:
$A=\frac{\lambda\delta\beta}{\alpha\mu\kappa}-\frac{\gamma}{\kappa}.$ (7)
This last result leads to the second stationary solution of the model, namely
$(R,A,P)=\left(\frac{\beta}{\alpha},\frac{\lambda\delta\beta}{\alpha\mu\kappa}-\frac{\gamma}{\kappa},\frac{\mu}{\lambda}\right).$ (8)
In analogy with critical phenomena, Eq. (7) can be rewritten in the form $A\sim(\lambda-\lambda_{c})^{\beta^{{}^{\prime}}}$, where the critical exponent is $\beta^{{}^{\prime}}=1$, and the critical defense threshold is given by:
@@
This result establishes a fundamental relationship between the size of the Roman army and the effectiveness of Syracuse’s defenses, represented by $\lambda$. If $\lambda$ exceeded the critical value given by Eq. (9), the city’s defenses could have successfully repelled the Roman forces. Conversely, since historically Syracuse fell, this suggests that $\lambda$ was below the critical threshold, confirming the overwhelming military pressure exerted by Rome.
The derived critical value $\lambda_{c}$ represents a key threshold in the siege dynamics. $\lambda<\lambda_{c}$, Syracuse would have been able to resist and potentially defeat the Roman forces. On the other hand, if $\lambda>\lambda_{c}$, Roman forces would ultimately overpower the defenders, leading to the historical outcome of Syracuse’s fall.
@@
And regarding the historical alignment? The actual historical scenario suggests that $\lambda>\lambda_{c}$, meaning Syracuse’s defenses, despite their ingenuity (including Archimedes’ war machines), were ultimately insufficient to counteract Roman military power, supply chains, and persistence. If Syracuse had access to greater supplies ($\beta$), suffered fewer losses ($\gamma$), or had a more effective defensive strategy ($\lambda$), its survival might have been possible. Conversely, Rome’s success was largely due to its ability to reinforce troops ($\mu$) and maintain pressure over the prolonged two-year siege.
@@
Determining the initial conditions for our model presents a challenge due to the lack of precise historical data regarding the population of Syracuse and the size of the Roman army at the time of the siege. Ancient sources, such as Polybius [5] and Livy [1], provide qualitative descriptions of the events but lack concrete figures. Plutarch’s Life of Marcellus [2] offers additional insights into the role of the Roman commander, but again, numerical data is scarce. Modern analyses, such as those by Goldsworthy [3] and Lazenby [16], attempt to reconstruct army sizes and logistical aspects of the war but often rely on estimations. Given these uncertainties, we consider reasonable approximations based on the available historical literature, allowing us to explore plausible scenarios for the siege. Thus, we considered as initial conditions $A(0)=35,000,P(0)=75,000$ and for $R(0)$ we choose a random numerical value, $R(0)=100.0$.
@@
In Fig. 2, panel (d), we varied $\delta$, that governs how effectively the available resources support the population of Syracuse. Higher values indicate that resources contribute more efficiently to sustaining the population, whereas lower values mean that even with sufficient resources, the population might still decline due to inefficiencies, distribution issues, or internal conflicts. For $\delta=0.0050$ (low resource effectiveness), even if resources ($R$) are available, they do not significantly help sustain the population ($P$). This could represent logistical problems, corruption, or spoilage of food supplies. The population declines faster, increasing the likelihood of an early collapse of Syracuse. For $\delta=0.010$ (moderate resource effectiveness - reference case), we have the baseline scenario, where resources effectively contribute to population maintenance, but only to a limited extent. The population decline is slowed, but if resources deplete ($\beta/\alpha$) reaches a critical level, Syracuse still falls within a realistic historical timeframe. For $\delta=0.020$ (high resource effectiveness), resources strongly sustain the population. As long as some resources remain, the population persists for a longer time. This scenario suggests that, had Syracuse been able to use its resources more efficiently (e.g., through better rationing, supply chains, or external aid), the siege could have lasted much longer. Thus, a lower $\delta$ means that even with available resources, the population struggles to survive, leading to an earlier fall. A moderate $\delta$ aligns with historical expectations, allowing resistance for a reasonable period before collapse. A higher $\delta$ suggests that with better logistics or supply management, Syracuse might have resisted for much longer, potentially changing the course of the siege.
@@
In Fig. 3, panel (b), we varied $\mu$, that directly influences the resilience of the Roman army ($A$). A higher value means that the Romans are capable of sustaining their military presence despite losses, while a lower value suggests difficulties in maintaining or replacing troops over time. For $\mu=0.00005$ (low reinforcement rate), the Roman army struggles to replenish itself. If Syracuse’s defenses ($\lambda$) are strong enough, there is a possibility that the Romans could be repelled over time. This case corresponds to a scenario where logistical difficulties, disease, desertions, or losses in battle significantly weaken the Roman presence. For $\mu=0.0001$ (baseline scenario), the Roman army is replenished at a moderate rate, which aligns with historical reality. The army does not grow indefinitely, but it also does not vanish due to attrition. This case represents a prolonged but ultimately successful siege, where the Romans maintain a persistent presence despite setbacks. For $\mu=0.002$ (high reinforcement rate), the Roman army is quickly replenished, making it nearly impossible for Syracuse’s defenders to deplete the attacking forces. Even if the city had more resources and stronger defenses, the overwhelming ability of Rome to bring in new troops ensures that the siege will succeed in a much shorter timeframe. This scenario would align with an alternate history where the Romans committed far more resources to the siege, leading to a swift and decisive conquest. Thus, a low $\mu$ suggests that if Syracuse had held out long enough, Rome might have abandoned the siege due to unsustainable losses. A moderate $\mu$ aligns with the historical outcome, where the Roman army was persistent but not invulnerable, leading to a gradual conquest. A high $\mu$ suggests that the fall of Syracuse was inevitable, as Rome’s ability to maintain a constant army would ensure eventual success, regardless of other factors.
@@ 4 Final Remarks
Particularly, the inclusion of the term $+\mu A$ in the equation for the attacking Roman forces introduces a crucial aspect of historical realism into the model. This term accounts for the replenishment of Roman troops during the siege, a well-documented feature of Roman military campaigns [1]. Without it, the model would assume a continuous depletion of the Roman forces, which does not accurately reflect their logistical capabilities and strategic persistence. By incorporating such term, the model recognizes that Rome had the resources to sustain its military presence for extended periods. This factor is consistent with historical accounts that describe Rome’s ability to recruit and deploy fresh troops when necessary, preventing significant attrition from weakening their offensive. The replenishment of forces creates a more asymmetrical struggle, where Syracuse’s defenses (modeled through $\lambda$) must not only inflict losses but do so at a rate that exceeds the reinforcements. If $\lambda>\lambda_{c}$, the critical defense threshold, then the Roman forces grow over time, making the eventual fall of the city more likely.

### v2

@@ Abstract
The Siege of Syracuse (214 - 212 BC) was a decisive event in the Second Punic War, leading to the city’s fall to Rome despite its formidable defenses, including the war machines devised by Archimedes. In this work, we propose a mathematical model to describe the dynamics of the siege, incorporating the depletion of resources, the decline of Syracuse’s population, and the persistence of the Roman army. Our analysis reveals the existence of a critical threshold $\lambda_{c}$, which determines the outcome of the siege. This threshold marks a phase transition: if the effectiveness of Syracuse’s defenses, represented by $\lambda$, had exceeded $\lambda_{c}$, the city could have withstood the Roman assault. However, since history records Syracuse’s fall, we conclude that $\lambda<\lambda_{c}$. This result provides a quantitative framework to understand the inevitability of the city’s conquest and demonstrates how mathematical modeling can offer new insights into historical military conflicts. We also explore different scenarios and assess the impact of key factors such as siege duration, supply constraints, and defensive capabilities. The results provide insights into how strategic elements influenced the eventual fall of Syracuse and demonstrate the applicability of mathematical modeling in historical military analysis.
@@ 1 Introduction
The siege of Syracuse (214 - 212 BC) was a pivotal event during the Second Punic War, in which the Roman Republic besieged and ultimately conquered the powerful Greek city-state of Syracuse, located in Sicily. This siege marked a crucial moment in Rome’s expansionist ambitions and solidified its control over the Mediterranean. Syracuse, previously an ally of Rome under King Hiero II, shifted its allegiance to Carthage after Hiero’s death, prompting a Roman military response under the command of Marcus Claudius Marcellus [1].
@@
Related approaches have also been proposed to model the historical dynamics of complex societies. For instance, Gunduz [16] developed a mathematical model to describe the rise and fall of empires, using a system of nonlinear differential equations to capture the evolution of political power over time. This type of framework reinforces the relevance of applying simplified mathematical models to analyze the structural dynamics behind major historical transitions, such as the fall of Syracuse.
@@
The fall of Syracuse can be understood in terms of a critical threshold in the effectiveness of its defenses. In the mathematical model developed in this work, we identify a critical value $\lambda_{c}$ that marks a transition between two possible outcomes: the survival of Syracuse or its eventual conquest by Rome. This behavior resembles a phase transition in physics, where small variations in parameters lead to qualitatively different macroscopic states. If the defensive effectiveness $\lambda$ had exceeded the critical threshold $\lambda_{c}$, Syracuse might have been able to resist the Roman siege indefinitely. However, since historical records indicate the city’s fall, we conclude that $\lambda<\lambda_{c}$, confirming that the defenses, despite their ingenuity - including the war machines attributed to Archimedes - were ultimately insufficient to prevent Roman victory. This approach provides a novel perspective on the siege, connecting historical events to universal principles of dynamical systems and critical phenomena.
Thus, in this work, we develop a system of three differential equations to model the interactions among available resources, the besieged population, and the Roman attacking force. These equations capture the depletion of resources over time, the decline in the defending population due to starvation and combat, and the attrition of the Roman army. By solving these equations numerically, we explore different scenarios and assess the conditions under which Syracuse ultimately fell.
@@ 2 Mathematical model
To analyze the dynamics of the Siege of Syracuse, we develop a system of three differential equations representing the interaction among available resources, the besieged population, and the attacking Roman forces. The model captures key factors such as resource depletion, attrition due to starvation and combat, and the effectiveness of defensive and offensive strategies.
@@
The first equation models the decline of resources due to consumption by the population and natural degradation over time. The second equation captures population losses due to natural causes, starvation, and direct military confrontations, but it presents a term that models the survivance due to limited resources. The third equation represents the attrition of the Roman army caused by resistance from the defenders, and Roman army reinforcement. It is important to note that the model does not include non-combat-related deaths of Roman soldiers, such as those resulting from disease, malnutrition, or exhaustion. This simplification was made to preserve analytical tractability and to focus on the primary mechanisms driving the outcome of the siege.
@@ 3 Results
From Eq. (3), we find two steady-state solutions: $A=0$ and $P=\mu/\lambda$. Using the steady-state value of $R$, the long-time limit of Eq. (2) gives us:
$P=\frac{\delta\beta}{\alpha(\gamma+\kappa A)}.$ (5)
For the case where $A=0$, we obtain $P=\frac{\delta\beta}{\alpha\gamma}$, leading to the first stationary solution:
$(R,A,P)=\left(\frac{\beta}{\alpha},0,\frac{\delta\beta}{\alpha\gamma}\right).$ (6)
On the other hand, if $A\neq 0$, we found $P=\frac{\mu}{\lambda}$. However, this expression does not represent a physical stationary solution for $P$, as it contradicts physical expectations. In fact, for increasing values of $\mu$ (Roman army reinforcement rate), we would expect a decrease in the population of Syracuse $P$, not an increase as this expression suggests. Similarly, increasing $\lambda$ (the effectiveness of Syracuse’s defenses) should enhance the survival of the population, implying an increase in $P$, not a decrease. The expression $P=\frac{\mu}{\lambda}$ merely identifies the condition under which the Roman army stabilizes (i.e., $dA/dt=0$), but it does not correspond to the actual steady-state value of $P$. The latter must be obtained from Eq. (5), and it depends on the equilibrium value of $A$, as well as on some of the model’s parameters.
However, the condition for the existence of a steady-state solution with $A\neq 0$ can be obtained through a stability analysis of the model equations (1) - (3) (see Appendix  [^1] The mathematical approach is grounded in standard dynamical systems theory [17, 18, 19], and reveals a critical transition associated with a change in the stability of the Roman army’s presence.). This analysis reveals the presence of a critical defense threshold, $\lambda_{c}$, given by:
@@
As shown in the Appendix, the stationary solution $A=0$ is stable only when $\lambda>\lambda_{c}$, i.e., when the effectiveness of Syracuse’s defenses is sufficiently strong to eliminate the Roman army. On the other hand, for $\lambda<\lambda_{c}$, the Syracuse’s defenses are too weak to resist the Roman assault, allowing the invading force to persist and eventually prevail.
The derived critical value $\lambda_{c}$ represents a key threshold in the siege dynamics. $\lambda>\lambda_{c}$, Syracuse would have been able to resist and potentially defeat the Roman forces. On the other hand, if $\lambda<\lambda_{c}$, Roman forces would ultimately overpower the defenders, leading to the historical outcome of Syracuse’s fall.
@@
And regarding the historical alignment? The actual historical scenario suggests that $\lambda<\lambda_{c}$, meaning Syracuse’s defenses, despite their ingenuity (including Archimedes’ war machines), were ultimately insufficient to counteract Roman military power, supply chains, and persistence. If Syracuse had access to greater supplies ($\beta$), suffered fewer losses ($\gamma$), or had a more effective defensive strategy ($\lambda$), its survival might have been possible. Conversely, Rome’s success was largely due to its ability to reinforce troops ($\mu$) and maintain pressure over the prolonged two-year siege.
@@
Determining the initial conditions for our model presents a challenge due to the lack of precise historical data regarding the population of Syracuse and the size of the Roman army at the time of the siege. Ancient sources, such as Polybius [5] and Livy [1], provide qualitative descriptions of the events but lack concrete figures. Plutarch’s Live book [2] offers additional insights into the role of the Roman commander, but again, numerical data is scarce. Modern analyses, such as those by Goldsworthy [3] and Lazenby [20], attempt to reconstruct army sizes and logistical aspects of the war but often rely on estimations. Given these uncertainties, we consider reasonable approximations based on the available historical literature, allowing us to explore plausible scenarios for the siege. Thus, we considered as initial conditions $A(0)=35,000,P(0)=75,000$ and for $R(0)$ we choose a random numerical value, $R(0)=100.0$.
@@
In Fig. 2, panel (d), we varied $\delta$, that governs how effectively the available resources support the population of Syracuse. Higher values indicate that resources contribute more efficiently to sustaining the population, whereas lower values mean that even with sufficient resources, the population might still decline due to inefficiencies, distribution issues, or internal conflicts. For $\delta=0.0050$ (low resource effectiveness), even if resources ($R$) are available, they do not significantly help sustain the population ($P$). This could represent logistical problems, corruption, or spoilage of food supplies. The population declines faster, increasing the likelihood of an early collapse of Syracuse. For $\delta=0.010$ (moderate resource effectiveness - reference case), we have the baseline scenario, where resources effectively contribute to population maintenance, but only to a limited extent. The population decline is slowed, but if resources deplete reaches a critical level, Syracuse still falls within a realistic historical timeframe. For $\delta=0.020$ (high resource effectiveness), resources strongly sustain the population. As long as some resources remain, the population persists for a longer time. This scenario suggests that, had Syracuse been able to use its resources more efficiently (e.g., through better rationing, supply chains, or external aid), the siege could have lasted much longer. Thus, a lower $\delta$ means that even with available resources, the population struggles to survive, leading to an earlier fall. A moderate $\delta$ aligns with historical expectations, allowing resistance for a reasonable period before collapse. A higher $\delta$ suggests that with better logistics or supply management, Syracuse might have resisted for much longer, potentially changing the course of the siege.
@@
In Fig. 3, panel (b), we varied $\mu$, that directly influences the resilience of the Roman army ($A$). A higher value means that the Romans are capable of sustaining their military presence despite losses, while a lower value suggests difficulties in maintaining or replacing troops over time. For $\mu=0.00005$ (low reinforcement rate), the Roman army struggles to replenish itself. If Syracuse’s defenses ($\lambda$) are strong enough, there is a possibility that the Romans could be repelled over time. This case corresponds to a scenario where logistical difficulties, disease, desertions, or losses in battle significantly weaken the Roman presence. For $\mu=0.0001$ (baseline scenario), the Roman army is replenished at a moderate rate, which aligns with historical reality. The army does not grow indefinitely, but it also does not vanish due to attrition. This case represents a prolonged but ultimately successful siege, where the Romans maintain a persistent presence despite setbacks. For $\mu=0.0002$ (high reinforcement rate), the Roman army is quickly replenished, making it nearly impossible for Syracuse’s defenders to deplete the attacking forces. Even if the city had more resources and stronger defenses, the overwhelming ability of Rome to bring in new troops ensures that the siege will succeed in a much shorter timeframe. This scenario would align with an alternate history where the Romans committed far more resources to the siege, leading to a swift and decisive conquest. Thus, a low $\mu$ suggests that if Syracuse had held out long enough, Rome might have abandoned the siege due to unsustainable losses. A moderate $\mu$ aligns with the historical outcome, where the Roman army was persistent but not invulnerable, leading to a gradual conquest. A high $\mu$ suggests that the fall of Syracuse was inevitable, as Rome’s ability to maintain a constant army would ensure eventual success, regardless of other factors.
@@ 4 Final Remarks
Particularly, the inclusion of the term $+\mu A$ in the equation for the attacking Roman forces introduces a crucial aspect of historical realism into the model. This term accounts for the replenishment of Roman troops during the siege, a well-documented feature of Roman military campaigns [1]. Without it, the model would assume a continuous depletion of the Roman forces, which does not accurately reflect their logistical capabilities and strategic persistence. By incorporating such term, the model recognizes that Rome had the resources to sustain its military presence for extended periods. This factor is consistent with historical accounts that describe Rome’s ability to recruit and deploy fresh troops when necessary, preventing significant attrition from weakening their offensive. The replenishment of forces creates a more asymmetrical struggle, where Syracuse’s defenses (modeled through $\lambda$) must not only inflict losses but do so at a rate that exceeds the reinforcements. If $\lambda<\lambda_{c}$, the critical defense threshold, then the Roman forces grow over time, making the eventual fall of the city more likely.
@@ Acknowledgments
Appendix A Stability analysis
The local stability of the model’s stationary solution can be inferred from the eigenvalues of the Jacobian matrix $J$ obtained from the system of differential equations (1) - (3). An equilibrium point is locally asymptotically stable if all eigenvalues of $J$ have negative real parts [17, 18, 19]. The eigenvalues can be obtained from $det(J-\Lambda\,I)=0$, where $I$ is the identity matrix. The Jacobian matrix is given by taking the partial derivatives of Eqs. (1) - (3), namely
$J=\begin{bmatrix}\frac{\partial{\dot{R}}}{\partial{R}}&\frac{\partial{\dot{R}}}{\partial{P}}&\frac{\partial{\dot{R}}}{\partial{A}}\\ \frac{\partial{\dot{P}}}{\partial{R}}&\frac{\partial{\dot{P}}}{\partial{P}}&\frac{\partial{\dot{P}}}{\partial{A}}\\ \frac{\partial{\dot{A}}}{\partial{R}}&\frac{\partial{\dot{A}}}{\partial{P}}&\frac{\partial{\dot{A}}}{\partial{A}}\end{bmatrix}$
Taking the derivatives, the matrix becomes
$J=\begin{bmatrix}-\alpha&0&0\\ \delta&-\gamma-\kappa\,A&-\kappa\,P\\ 0&-\lambda\,A&\mu-\lambda\,P\end{bmatrix}$
For the solution $(R,A,P)=(\beta/\alpha,0,\delta\beta/\alpha\gamma)$, the eigenvelues of the Jacobian matrix are $\Lambda_{1}=-\alpha,\Lambda_{2}=-\gamma$ and $\lambda_{3}=\mu-\lambda\delta\beta/\alpha\delta$. We have $\Lambda_{1}<0$ and $\Lambda_{2}<0$ since $\alpha>0$ and $\gamma>0$. We have $\Lambda_{3}<0$ only if
$\lambda>\frac{\alpha\gamma\mu}{\delta\beta}$ (8)
From Eq. (8), we identify
$\lambda_{c}=\frac{\alpha\gamma\mu}{\delta\beta}$ (9)
In other words, the solution $A=0$ (extinction of the Roman army) is stable only if $\lambda>\lambda_{c}$, which is the condition for the survival of Syracuse. Conversely, if $\lambda<\lambda_{c}$, the Roman army persists, ultimately leading to Syracuse’s fall (which is related to a solution $A\neq 0$).